how do you find all possible solutiions of 3tanx * cosx = -5cosx?
Answer by user_dude2008(1862)

3tanx * cosx = -5cosx


3tanx * cosx + 5cosx = 0

cosx(3tanx + 5) = 0


cosx = 0 or 3tanx + 5 = 0

x = pi/2+2pi*n or x = -1.03037 + 2pi*n